OpenStudy (anonymous):

Solve the equation. –6 + 3x = –9 A. –1 B. –6 C. –5 D. –3

OpenStudy (king):

3x=-3 x=-1

OpenStudy (bahrom7893):

-6+3x=-9 3x=-9+6 3x=-3 x=-1